### v1

# Controllable non-Hermitian qubit-qubit Coupling in Superconducting quantum Circuit

Abstract
With a high-loss resonator supplying the non-Hermiticity, we study the Energy level degeneracy and quantum state evolution in tunable coupling superconducting quantum circuit. The qubit’s effective energy level and damping rate can be continually tuned in superconducting circuit, and the positions and numbers of level degenerate points are controllable. The efficient of quantum state exchange and the asymmetry of quantum state evolution can be tuned with non-hermitian and nonreciprocal coupling between two qubits. The controllable non-Hermiticity provides new insights and methods for exploring the unconventional quantum effects in superconducting quantum circuit.
I Introduction
The non-Hermitian system possessing complex energy spectrum, the real and imaginary parts respectively label the energy levels and loss of system, while the degenerate point of energy levels are known as the Exceptional points[3, 4, 5, 6, 7]. The system close to exceptional points contain rich unconventional physical effects, including the ultralow driving threshold chaos[23], high sensitivity metrology[9, 10], high-dimensions skin effect[11, 12], Chiral Heat Transport[13] and so on. The exceptional point have been studied in single ion system[14, 15], optical microcavity[16, 17, 18, 19, 20, 21, 22, 24], and optomechanical system[25, 26, 27, 28], magnonic system[29, 30, 31], Hybrid Quantum System[23], single superconducting qubit[32, 33], and so on. Some theoretical work have studied the PT-symmetry in superconducting circuit with the single qubit coupling to the resonatorcite[34, 35]. However, it is still challenging to experimentally observe such non-Hermitian in multi-qubits superconducting circuit due to the difficulties in realizing controlled gain and loss.
Unlike the efforts to improve the gain and loss of superconducting resonator, we use a high-loss resonator supplying the effective non-Hermitian interactions between two qubits in tunable coupling superconducting circuit. Using the advantage of easy tuning for superconducting artificial atoms, the energy level degenerate and Exceptional points of the non-Hermitian superconducting circuit are controlled by the phases and amplitudes of qubit-resonator coupling, the effective qubit-qubit coupling, and the quantum state of qubits. Because of special parameter dependent relation in tunable coupling superconducting circuit, we find that the energy level degenerate point can be one or a series in different parameter regimes, and the positions of the exceptional points are controllable. The forward and backward coupling between two qubits are still non-reciprocal in non-Hermitian Superconducting circuit, which can be used to develop the integrated unidirectional transmission devices.
By preparing one qubit in the excited state and another in the ground states, the free evolution of quantum states under non-hermite coupling can be tuned with the J-C coupling phases. Even setting the identical paramters, the evolutions of two qubits’ quantum states are different from different coupling directions. The non-reciprocal couplings between in superconducting circuit can be used to develop the integrated unidirectional transmission devices including the microwave circulator and isolator.
The paper is organized as follows: In Sec. II, we build the model of non-Hermitian superconducting circuit; In Sec. III, the energy level degeneracy and the quantum state evolution are tuned in energy level attraction regimes; In Sec. IV, we analyze the non-reciprocal coupling and asymmetric quantum state evolutions. We finally summarize the results in Sec. V.
II Effective Non-Hermitian Circuit
As shown in Fig. 1, we study a superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss superconducting resonator. The transmon coupler and the high-loss resonator can induce the hermite and non-hermite indirect interaction between two qubits. Defining $\omega_{k}$ as the transition frequencies of superconducting artificial atoms, and $\sigma^{(k)}_{z}$ and $\sigma^{(k)}_{\pm}$ are the respective pauli-Z and ladder operators, where $k=1,2,c$ label the qubit 1, qubit 2 and transmon coupler, respectively.
The free Hamiltonian for the superconducting circuit can be written as $H_{0}=\omega_{a}a^{\dagger}a+\sum_{k=1,2,c}(\omega_{k}/2)\sigma^{(k)}_{z}$. $\omega_{a}$ is the resonant frequency of resonator, and the corresponding transition frequency of superconducting artificial atoms are defined as $\omega_{k}$, where $k=1,2,c$ label the qubit 1, qubit 2, and the transmon coupler, respectively. $a^{\dagger}$ and $a$ are the creation and annihilation operators for the resonator photons, and $\sigma^{(k)}_{z}$ is the Pauli-Z operator of superconducting artificial atom. The Hamiltonian for the direct qubit-qubit interaction is $H_{xy}=g_{xy}[\sigma^{(1)}_{+}\sigma^{(2)}_{-}+\sigma^{(2)}_{+}\sigma^{(1)}_{-}]$, where $g_{xy}$ label the direct coupling strength between two qubits. We defined the $\sigma^{(k)}_{\pm}$ as the ladder operators of artificial atoms. The Hamiltonian for the interaction between coupler and qubit-$j$ can be written as $H^{j}_{qc}=g_{j}[\sigma^{(c)}_{+}\sigma^{(j)}_{-}+\sigma^{(j)}_{+}\sigma^{(c)}_{-}]$, where $g_{j}$ is the coupling strength between qubit-$j$ and transmon coupler, with $j=1,2$.
Figure 1: (Color online) The superconducting circuit consisting of two Xmon qubits coupling to a common transmon coupler and a high-loss resonator. The transition frequency of superconducting artificial atom is $\omega_{k}$, where $k=1,2,c$ label the qubit 1, qubit 2, and transmon coupler, the resonant frequency of resonator is $\omega_{a}$. The capacitance type interaction is dominated between arbitrary two devices.
In this article, the frequency detuning microwave is applied on qubit to create amplitude and phase tunable J-C (Jaynes Cummings) interaction between qubit and resonator[36, 37, 38]. The qubits and resonator are in the dispersive coupling regimes, that is $\lambda^{(0)}_{j}\ll|\Delta_{ja}|,|\Delta_{ja}+\alpha_{j}|$, where $\lambda^{(0)}_{j}$ labels the bare J-C coupling amplitude between first-excited state of superconducting artificial atom-$j$ and the fundamental mode of resonator and their frequency detuning is $\Delta_{ja}=\omega_{j}-\omega_{a}$, while $\alpha_{j}$ describe the anharmoncity of qubit-$j$. After injecting a single tone microwave pulse with amplitude $\Omega_{j}(t)$ and phase $\theta_{j}(t)$, thus J-C interaction between resonator and qubit-$j$ can be obtained as $\lambda_{j}\exp{(\theta_{j})}$, with $\lambda_{j}=(1/\sqrt{2})\lambda^{(0)}_{j}\alpha_{j}|\Omega_{j}|/[\Delta_{ja}(\Delta_{ja}+\alpha_{j})]$[36, 37]. The amplitude $\lambda_{j}$ and phase $\theta_{j}$ can be tuned by the input microwave pulse, thus the Hamiltonian for amplitude and phase tunable J-C coupling can be written as $H^{j}_{qr}=[\lambda_{j}\exp{(-i\theta_{j})}a^{\dagger}\sigma^{(j)}_{-}+\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{+}a]$. The interaction between resonator and transmon coupler is very weak and can be neglected, thus the Hamiltonian for the superconducting circuit in Fig. 1 can be obtained as $H=H_{0}+H_{xy}+\sum_{j}H^{j}_{qc}+\sum_{j}H^{j}_{qr}$.
In the weak dispersive coupling regimes $g_{j}/|\Delta_{jc}|\sim 1/3$ (with $\Delta_{jc}=\omega_{j}-\omega_{c}$), the Schrieffer-Wolff transformation $U=\exp{\sum_{j=1,2}(g_{j}/\Delta_{jc})[\sigma^{(c)}_{+}\sigma^{(j)}_{-}-\sigma^{(j)}_{+}\sigma^{(c)}_{-}]}$ can be applied to Hamiltonian $H$ to cancel the qubit-coupler interaction terms[39, 40]. Thus the effective Hamiltonian of two qubits coupling to a common high-loss resonator can be written as
$\frac{H^{\prime}}{\hbar}$ $=$ $\omega_{a}a^{\dagger}a+\left(\omega_{j}+\sum_{j=1,2}\frac{g^{2}_{j}}{\Delta_{jc}}\right)\frac{\sigma^{(j)}_{z}}{2}$
$+$ $g_{e}\left[\sigma^{(1)}_{+}\sigma^{(2)}_{-}+\sigma^{(2)}_{+}\sigma^{(1)}_{-}\right]$
$+$ $\sum_{j=1,2}\lambda_{j}\bigg{[}\exp{(-i\theta_{j})}a^{\dagger}\sigma^{(j)}_{-}+\exp{(i\theta_{j})}\sigma^{(j)}_{+}a\bigg{]},$
where the subscripts $j=1,2$ label the qubit 1 and qubit 2, respectively. The effective interaction between two qubits is defined as $g_{e}=g_{xy}+g_{1}g_{2}/\Delta_{e}$ with $2/\Delta_{e}=1/\Delta_{1c}+1/\Delta_{2c}$, where the first term is the direct qubit-qubit interaction and the second term labels the indirect interaction induced by the transmon coupler[41, 42, 43]. The indirect qubit-qubit interaction induced by the high-loss resonator is the non-hermite type and not contained in $g_{e}$. Since $g_{xy}\ll g_{j}$, the invariance assumption is made for $H_{xy}$ during the Schrieffer-Wolff transformation, and the tunable coupler is assumed as the ground states and thus the free Hamiltonian of transmon coupler is neglected.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: (Color online) Controllable levels degeneracy in non-hermitian superconducting circuit. (a) Effective energy levels and (b) damping rates as the function of effective coupling $g_{e}$: $\Delta\theta=0.500\pi$ (blue curves), $0.501\pi$ (red curves), and $0.502\pi$ (black curves). The two qubits are assumed in the ground states $\langle\sigma^{j}_{z}\rangle=-1$, with $j=1,2$. (c) Effective energy levels as the function of J-C coupling amplitude $\lambda_{j}$: $\Delta\theta=0.500\pi$, $\Gamma_{2}=\Gamma_{1}$, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz (blue curves); $g_{e}/2\pi=1.224$ MHz (green curves); $\gamma_{a}/2\pi=66$ MHz (red curves); $\Gamma_{2}=0.99\Gamma_{1}$(black curves). $\langle\sigma^{(1)}_{z}\rangle=-1$, and other un-mentioned parameters on the green, red, and black curves are the same as that in blue curves. (d) Effective energy levels as function of qubit populations: $\Delta\theta=0.500\pi$, $\lambda_{j}/2\pi=11$ MHz, $g_{e}/2\pi=1.137$ MHz (blue curves); $\Delta\theta=0.5005\pi$, the other parameters same as blue curve (red curves); $g_{e}/2\pi=1.224$ MHz, $\lambda_{j}/2\pi=11$ MHz, other parameters same as blue curves (black curves). The population of two qubits are assumed oppositive, $\langle\sigma^{j}_{z}\rangle=(-1)^{(j+1)}\langle\sigma_{z}\rangle$) The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
In the rotating coordinates at resonator’s frequency $\omega_{a}$, the motion equations of resonator and qubits can be written as $\dot{a}=-\gamma_{a}a-i\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$ and $\dot{\sigma}^{(j)}_{-}=-\left[\gamma_{j}/2+i\left(\Delta_{ja}+g^{2}_{j}/\Delta_{jc}\right)\right]\sigma^{(j)}_{-}-i\lambda_{j}\exp{(i\theta_{j})}\sigma^{(j)}_{z}a-ig_{e}\sigma^{(j)}_{z}\sigma^{(j+1)}_{-}$, with $\Delta_{ja}=\omega_{j}-\omega_{a}$ and $j+1=1$ for $j=2$. In the case of $\gamma_{a}\gg g_{j},\lambda_{j}\gg\gamma_{j}$ and $|\Delta_{ja}|/\gamma_{a}\approx 1/3$, with the adiabatic approximation, we can get $a=-i(\lambda_{1}/\gamma_{a})\sum_{j=1,2}\lambda_{j}\exp{(-i\theta_{j})}\sigma^{(j)}_{-}$. Substituting it into the motion equation of ladder operators, thus we get
$\dot{\sigma}^{(j)}_{-}$ $=$ $-\left[\left(\frac{\gamma_{j}}{2}+\frac{\lambda^{2}_{j}}{\gamma_{a}}\sigma^{(j)}_{z}\right)+i\left(\Delta_{ja}+\frac{g^{2}_{j}}{\Delta_{jc}}\right)\right]\sigma^{(j)}_{-}$ (2)
-- $\left[ig_{e}+\frac{\lambda_{j}\lambda_{j+1}}{\gamma_{a}}\exp{[i(\theta_{j}-\theta_{j+1})]}\right]\sigma^{(j)}_{z}\sigma^{(j+1)}_{-}.\quad$
The Pauli-Z operator $\sigma^{(j)}_{z}$ describes the population of qubit-$j$ and its steady value can be tuned with the external microwave signals. Though coupling to a high-loss resonator, the decoherence time of qubit is still much longer than the durations for preparation and readout of single qubit’s quantum state (usually below 100 ns) especially for the tantalum-based superconducting qubit[44, 45].
Thus the population of qubit-$j$ can be considered as staying in the transient stability value $\langle\sigma^{(j)}_{z}\rangle$ during the preparations and measurements of qubits’ quantum states, and its value can be tuned by the qubit-coupler coupling strength, and also the J-C coupling phases and amplitudes. Thus the effective Hamiltonian for non-Hermitian qubit-qubit coupling can be written as
$\frac{H^{\prime\prime}}{\hbar}=\left(\begin{array}[]{cc}\frac{1}{2}(\Delta^{\prime}_{1a}-i\Gamma_{1})&g_{e}-\frac{i\lambda_{1}\lambda_{2}}{\gamma_{a}}\exp{(i\Delta\theta)}\\ g_{e}-\frac{i\lambda_{1}\lambda_{2}}{\gamma_{a}}\exp{(-i\Delta\theta)}&\frac{1}{2}(\Delta^{\prime}_{2a}-i\Gamma_{2})\\ \end{array}\right),$ (5)
where $\Delta\theta=\theta_{1}-\theta_{2}$, $\Delta^{\prime}_{ja}=\Delta_{ja}+g^{2}_{j}/\Delta_{jc}$, and $\Gamma_{j}=\gamma_{j}-\lambda^{2}_{j}\langle\sigma^{(j)}_{z}\rangle/\gamma_{a}$, with $j=1,2$. The non-diagonal terms in Eq.(3) label total effective qubit-qubit coupling strength in the nonhermite superconducting circuit. The two Eigen-modes for the non-Hermitian Hamiltonian $H^{\prime\prime}$ can be calculated as
$\omega_{\pm}=\frac{\Delta^{\prime}_{1a}+\Delta^{\prime}_{2a}}{4}-i\frac{\Gamma_{1}+\Gamma_{2}}{4}\pm\sqrt{R+iI},$ (6)
where we have defined
$R$ $=$ $\frac{(\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a})^{2}}{4}-\frac{(\Gamma_{1}-\Gamma_{2})^{2}}{4}+4g^{2}_{e}-\frac{4\lambda^{2}_{1}\lambda^{2}_{2}}{\gamma^{2}_{a}},$ (7)
$I$ $=$ $-\frac{8g_{e}\lambda_{1}\lambda_{2}}{\gamma_{a}}\cos(\Delta\theta)-\frac{(\Gamma_{1}-\Gamma_{2})(\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a})}{2}.$ (8)
The real part $\Re{(\omega_{\pm})}$ corresponds to the effective energy levels of two qubits, while the imaginary part $\Im{(\omega_{\pm})}$ describes effective damping rates. The values of $R$ and $I$ affect the separations of complex energy spectrum between two Eigen-modes, and the signs of $R$ decide the level repulsions or attractions, while the sign of $I$ describes the relative size of two effective damping rates.
III Controllable Non-Hermitian coupling
The experimental tunable parameters in superconducting circuit of Fig. 1 include the J-C coupling phases ($\theta_{j}$) and amplitudes $\lambda_{j}$, effective qubit-qubit coupling $g_{e}$, and population $\sigma^{(j)}_{z}$ of qubit $j$, which can be continually tuned to find the available parameters for exceptional point. The $\sqrt{R+iI}$ becomes a pure imaginary number($\Delta E_{q}=0$) in the case of $R<0$ and $I=0$, so the energy levels of two qubits can merge as one and the exceptional point should appear. The two qubits will possess misaligned energy levels even for a small nonzero value of $I$, and the exceptional point should be absent[28].
By setting $\langle\sigma^{(j)}_{z}\rangle=-1$ ($\Gamma_{1}-\Gamma_{2}\neq 0$) and J-C coupling amplitude $\lambda_{j}/(2\pi)=11$ MHz, the effects of J-C coupling phases on the effective energy levels and damping rates of two Eigen-modes are shown in Figs. 2(a) and  2(b), respectively. The separations of two qubits’ effective energy levels (or damping rates) becomes narrower (or wider) in certain parameter regimes corresponding to the the energy level attraction. To see clearer, the curves in blue and black color shift up -0.5 MHz or 0.5 MHz in Fig. 2(a), respectively. As shown in the inset figure of Fig. 2(a), there is a degenerate points on the pairs of two red or black color curves for $\Delta\theta\neq\pi/2$. In the case of $\Delta\theta=\pi/2$, the first term in the right side of Eq.(6) is always zero, since $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in parameter area of drawing, so there is no level degenerate point on the blue color curves in Fig. 2(a). The scattering curves for damping rates of two Eigen-mode are shown in Fig. 2(b), to see clearer, the curves in red and black color shift up 0.1 MHz and 0.2 MHz, respectively. The two branch of effective damping rates exchange with each other in two blue or black color curves in Fig. 2(b), and the corresponding exchange positions coincides well with the corresponding locations of level degenerate points in Fig. 2(a), which should indicate the positions of Exceptional points.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: (Color online) Controllable quantum state evolutions in energy level attraction regimes. (a) Differences in effective energy level ($\Delta E_{q}$) and (b) in effective damping rate ($\Delta\gamma_{q}$) as the functions of phase difference $\Delta\theta$ and coupling strength $g_{e}$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The free evolutions for excited state occupation populations for qubits (c) starting in excited state or (d) ground state for the phase differences: $\Delta\theta=0$ (blue-solid curves); $\Delta\theta=\pi/3$ (green-dashed curves); $\Delta\theta=5\pi/12$ (red-dotted curves); $\Delta\theta=\pi/2$ (black-dash-dotted curves). A small effective coupling ($g_{e}/2\pi=-80$) kHz to realize the energy level attractions, and the other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, and $\gamma_{a}/2\pi=65$ MHz.
The effects of J-C coupling amplitudes on energy levels of two qubit are studied in Fig. 2(c). In the case of $\Delta\theta=\pi/2$ and $\Gamma_{1}-\Gamma_{2}=0$, the first and second terms in Eq.(6) are always zeroes, so the energy level degeneracy appear for arbitrary value of $\lambda_{j}$ (energy level attraction regimes) in the blue curves. The change for the values of $g_{e}$ can shift the position of the Exception point and also the value of energy levels as indicated by the green curves (compared with the blue curves). If the $\Gamma_{1}-\Gamma_{2}\neq 0$, because of no solution for $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in the parameter area of drawing, so the energy level degenerate points does not appear in the red and black curves. As indicated by the motion equation of $\sigma^{(j)}_{z}$, the effective energy levels of two qubits can be tuned by changing the quantum states of qubits.
The value of $\Gamma_{1}-\Gamma_{2}$ depends on the populations $\langle\sigma^{(j)}_{z}\rangle$ of qubit-$j$, and the energy level degenerate point can appear when the quantum states of qubits satisfy $\Gamma_{1}-\Gamma_{2}=0$ as indicated by the blue ($\Delta\theta=\pi/2$) and red-dashed ($\Delta\theta=0.5005\pi$) curves. Since the $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}\neq 0$ in the parameter area of drawing and the first term of Eq.(6) is zero in the case of $\Delta\theta=\pi/2$, so the $I=0$ can not be satisfied in the parameter regimes of $\Gamma_{1}-\Gamma_{2}\neq 0$ and there is no energy level degenerate points in the black-dashed curve. To change the transient value of $\langle\sigma^{(j)}_{z}\rangle$, the qubit-$j$ should be firstly initialized to the ground states and then excited to a new quantum states with the microwave pulses.
The separations between two effective energy levels and two damping rates of Eigen-modes can be defined as $\Delta E_{q}=\Re{\omega_{+}}-\Re{\omega_{-}}=2\Re{(\sqrt{R+iI})}$ and $\Delta\gamma_{q}=\Im{\omega_{+}}-\Im{\omega_{-}}\approx 2\Im{(\sqrt{R+iI})}$, respectively. In the case of $\langle\sigma^{(j)}_{z}\rangle=-1$ ($\Gamma_{1}-\Gamma_{2}\neq 0$), the differences in energy levels ($\Delta E_{q}$) and in damping rates ($\Delta\gamma_{q}$) as the function of effective coupling strength $g_{e}$ and phase difference $\Delta\theta$ are plotted in Figs. 3(a) and 3(b), respectively. In the parameter regimes of small values for $\Delta^{\prime}_{1a}-\Delta^{\prime}_{2a}$, $\Gamma_{1}-\Gamma_{2}$, and $\lambda_{1}\lambda_{2}/\gamma^{2}_{a}$, the value of $R$ is smaller for weaker effective coupling $g_{e}$ according to Eq.(5), and the value of $I$ is close to zero in the case of $\cos{\Delta\theta}\approx 0$. Thus the levels difference $\Delta E_{q}$ gets local minimal values close to the regimes of $g_{e}\rightarrow 0$ or $\Delta\theta\rightarrow\pi/2$ in Fig. 3(a). The sign of $I$ decides the relative size of two effective damping rates, thus the sign of $\Delta\gamma_{q}$ changes between two sides of the line $g_{e}\approx 0$ or $\Delta\theta\approx\pi/2$ in Fig. 3(b). The exceptional points can be found at the parameter regimes close $\Delta E_{q}\rightarrow 0$ and $\Delta\gamma_{q}\rightarrow 0$ in Figs. 3(a) and  3(b).
The free evolutions of quantum states under non-hermitian coupling can be calculated with the semiclassical limit method by disregarding the quantum jump terms[33, 49, 50]. By preparing the qubit 1 in the excited states and qubit 2 in the ground states, the time evolution of occupation probabilities for qubits’ excited state can be calculated with the semiclassical master equation $\dot{\rho}=-i(H^{\prime\prime}\rho-\rho H^{\prime\prime})$, where the $\rho$ is the density operator. To keep the superconducting circuit in the level attraction regimes, the effective qubit-qubit coupling $g_{e}$ is tuned to a small value, $g_{e}/2\pi=-80$ kHz. The time evolution of occupation probabilities for qubit-$j$ starting in excited ($P^{(se)}_{j}$) or ground ($P^{(sg)}_{j}$) state is shown in Fig. 3(c) or Fig. 3(d), respectively. The periodic oscillation of the curves originates from the nonzero frequency detuning $\Delta_{ja}$ in $H^{\prime\prime}$, and the effects of the J-C coupling phase can be reflected by the envelopes of the oscillating curves.
The larger oscillating amplitudes in the black-dash-dotted curves for $\Delta\theta\approx\pi/2$ originates from the small damping rate contributed by non-hermite coupling close to exceptional points. The maximal amplitude of non-hermite coupling strength is defined as $\Omega_{n}=\lambda_{1}\lambda_{2}/\gamma_{a}$, the positions of local maximal or minimal values for the envelope are different on the four curves which originate from the dependencies of total qubit-qubit coupling on the J-C coupling phases. The different values of local maximal values on envelope imply that the efficient of quantum state exchange between two qubits can be tuned by the J-C coupling phase, and the exchange efficient gets the highest value close to the exception point. In the black curves ($\Delta\theta=\pi/2$), the envelope of occupation probability drops slowly compared with the results on the blue-solid ($\Delta\theta=0$), green-dashed ($\Delta\theta=\pi/3$), and red dash-dotted ($\Delta\theta=5\pi/12$) curves, which indicates the different contributions of the non-hermite coupling on the effective damping rates of qubits. The quantum state exchange efficient and the costing time getting local maximal values can be tuned with the J-C coupling phases which might be useful to enhance the speed and fidelity of two-qubit quantum gates.
IV nonreciprocal coupling in Non-Hermitian circuit
As indicted by the previous work, the non-reciprocal transmission can be realized in non-Hermitian Circuit[46, 47, 48]. According to Eq.(3), the non-Hermitian coupling strengths between two qubits in the forward and backward directions can be respectively defined as $g^{(q)}_{1\rightleftharpoons 2}=g_{e}-i\Omega_{n}\exp{(\mp i\Delta\theta)}$, where the first term $g_{e}$ is hermite interactions, while the second-term $i\Omega_{n}\exp{(\mp i\Delta\theta)}$ is the non-hermite qubit-qubit coupling induced by the high-loss resonator.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: (Color online) Controllable Non-reciprocal couplings and state evolution in non-Hermitian superconducting circuit. (a) Non-Hermitian coupling strengths $g^{(q)}_{1\rightarrow 2}$ and (b) $g^{(q)}_{1\leftarrow 2}$ as functions of phase difference ($\Delta\theta$) and effective coupling strength ($g_{e}$). The one-dimensional curves of $g^{(q)}_{1\rightarrow 2}$ (blue-dashed curve) and $g^{(q)}_{1\leftarrow 2}$ (red-solid curve) for (c) $\Delta\theta=\pi/3$ and (e) $\Delta\theta=\pi/2$. The one-dimensional curves of non-reciprocal ratio $|g^{(q)}_{1\leftarrow 2}|/|g^{(q)}_{1\rightarrow 2}|$ for (d) $\Delta\theta=\pi/2$ and (f) $\Delta\theta=\pi/3$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz. The differences in occupation probabilities for two qubits starting in (g) excited states ($P^{(e)}_{j}$) and (h) ground ($P^{(g)}_{j}$) states. The two qubits adopt same parameters:$\omega_{j}/2\pi=4.5$ GHz, $g_{j}/2\pi=30$ MHz, and $\gamma_{j}/2\pi=1$ MHz (with $j=1,2$), and the other parameters are the same as in (a)-(f) except for $g_{e}/2\pi\approx-80$ KHz.
The $g^{(q)}_{1\rightarrow 2}\neq g^{(q)}_{1\leftarrow 2}$ even in the case of $\Delta\theta=n\pi$ and the local maximal (or minimal) values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ are out of steps as shown in Figs. 4(a) and 4(b), which implies the non-reciprocal coupling for two different directions of effective qubit-qubit coupling. The minimal values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ could reach zero in Fig. 4(e) for $\Delta\theta=\pi/2$, but they get nonzero minimal value in the case of $\Delta\theta=\pi/3$ in Fig. 4(c). And the nonreciprocal ratio is much stronger in the case of $\Delta\theta=\pi/2$ as indicated by the maximal and minimal values of $|g_{1\leftarrow 2}|/|g_{1\rightarrow 2}|$ in Figs. 4(d) and  4(f). The non-reciprocal coupling strengths between two qubits can also be tuned by the J-C coupling amplitudes and qubits’ quantum states.
The nonreciprocal coupling between two qubits might also result in the asymmetry of quantum state evolutions from different directions, which can be seen from the nonzero values for difference in occupation probabilities of two qubits’ excited states (same initial states) in Figs. 4(g) and  4(h). Beside the apparently asymmetric of occupation probabilities after the period $t>1/\Omega_{n}$, the nonreciprocity of qubit quantum state evolution can also be tuned by the J-C coupling phases. Even adopting the same frequencies, same coupling strengths, same damping rates, and same initial states, the time evolution of qubit occupation probabilities from two different directions still do not overlap in Figs. 4(g) and  4(h), which coincides with the results of non-reciprocal coupling in Figs. 4(c) and  4(f).
V Conclusions
In conclusion, the energy level degeneracies and quantum state evolutions are studied in non-Hermitian superconducting circuit. The non-hermite and nonreciprocal coupling between two qubits also affect the evolution of qubits’ quantum states, the efficient and asymmetry of quantum state exchange can be tuned by the J-C coupling phase which might be useful for the quantum gates. The resonator and qubits are in weak dispersive coupling regimes, so the state leakages and crosstalk are inevitable which should be suppressed with tunable coupler and wave-shape correction techniques[42].
VI ACKNOWLEDGMENTS
We thanks Zhiguang Yan for useful discussions.. H.W. is supported by the Natural Science Foundation of Shandong Province under Grant No. ZR2023LZH002 and the Inspur artificial intelligence research institute. Y.J.Z. is supported by Beijing Natural Science Foundation under Grant No. 4222064 and NSFC under Grant No. 11904013. X.-W.X. is supported by the National Natural Science Foundation of China (Grants No. 12064010 and No. 12247105), the Science and Technology Innovation Program of Hunan Province (Grant No. 2022RC1203), Natural Science Foundation of Hunan Province of China (Grant No. 2021JJ20036), and Hunan provincial major sci-tech program (Grant No. 2023ZJ1010).